OpenStudy (laronjaxon):

Solve x2 + 10x + 24 = 0 by completing the square. Show your work for full credit.

OpenStudy (lalaly):

\[x^2+10x+24=0\]\[(x+\frac{10}{2})^2-(\frac{10}{2})^2+24=0\]\[(x+5)^2-25+24=0\]\[(x+5)^2-1=0\]

OpenStudy (anonymous):

x^2+10x+24+1-1=0 (x+5)^2-1=0 (x+6)(x+4)=0

OpenStudy (lalaly):

\[(x+5)^2=1\]\[x+5= \pm \sqrt{1}\]\[x=-5 \pm \sqrt{1}\]